# A General Theory of Abiogenesis: Thermodynamic Selection and the Emergence of Life

### v5

@@ 3 Core Thermodynamic Framework
4 Modulators of Persistence
Entropy, Diffusion, and Spatial Organization
While entropy is traditionally framed as a measure of disorder, in the context of open thermodynamic systems, it plays a more nuanced and generative role. Systems far from equilibrium, such as prebiotic chemical environments, can increase global entropy while locally producing order through self-organization and energy dissipation (Schneider & Sagan, 2005; Prigogine and Nicolis, 1977).
To visualize this, imagine entropy as a crowd entering a large hall. Left unconstrained, people spread to fill the room evenly. But if pathways, walls, or rails guide their movement, order emerges, not by suppressing motion, but by channeling it. Entropy operates similarly in chemical systems: it enables exploration of configurations, while environmental structure filters for persistence.
We propose that it is precisely this dynamic, where entropy production coexists with local structure formation, that enables the emergence of chemically diverse distributions capable of sustaining consistent reactions across fluctuating energy input. In this view, entropy is not merely a passive constraint but a fundamental driver of molecular complexity.
These entropic transformations, guided by energy flux, yield a rich combinatorial space of molecular configurations. Among these, only subsets that align energetically with environmental fluctuation will persist, thereby introducing a thermodynamic basis for selection without heredity.
Thus, entropy facilitates the exploration of chemical state space, while energy-based viability filters that space to favor persistent configurations. This interplay underpins the reaction viability inequality and reframes entropy as a formative precursor to natural selection itself.
Similarly, diffusion and spatial factors constrain the thermodynamic landscape. Reaction viability depends not only on energetic balance but also on spatial proximity of reactants, molecular crowding, and the presence of gradients or compartments that limit dispersion. Structured environments, such as mineral surfaces, vesicles, or pores, may facilitate energy retention and constrain entropy in ways that support persistence.
Extension of the Inequality with Entropic Factors
To preserve the simplicity and elegance of the original reaction viability inequality, we propose that entropy and spatial factors be treated as modulators of system persistence, influencing the feasibility of local order and structural coherence under fluctuating energy (Schneider & Sagan, 2005; Morowitz, 1968).
We introduce the term $\Delta S_{net}$ to represent the net entropy balance between the system and its environment:
$\Delta S_{net}=\Delta S_{env}-\Delta S_{local}$
where:
  • $\Delta S_{env}$ is the entropy exported to the surroundings (e.g., via heat, dispersal)
  • $\Delta S_{local}$ is the entropy reduction due to local self-organization or structure (Prigogine and Nicolis, 1977; Deamer, 2011)
For persistence to remain thermodynamically viable, $\Delta S_{net}$ must remain positive, meaning that localized decreases in entropy (e.g., formation of stable structures) are compensated by environmental entropy production (Schneider & Sagan, 2005). This relationship allows ordered reaction networks to form and persist without violating the second law.
Additionally, a spatial modifier $D(x,t)$ can be introduced to represent diffusion constraints, with low diffusion (e.g., due to compartmentalization or surface adherence) enhancing the retention of energy and structure (Rastogi, Kumar & Patnaik, 2022; Deamer & Dworkin, 2002).
To account for both entropy and diffusion constraints, we define an augmented viability function $y^{\prime}(t)$ as:
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t)$
where:
  • $y(t)$ is the original residual energy as defined earlier
  • $R_{n}$ is a resilience factor representing structural redundancy or robustness
  • $\Phi(t)$ is an entropic-diffusive penalty defined as:
    $\Phi(t)=f(\Delta S_{net}^{-1},D(x,t))$
The updated persistence condition becomes:
$y^{\prime}(t)\geq 0$
These considerations do not negate the core reaction viability inequality, but rather refine its application with physically meaningful constraints that can be applied when modeling realistic prebiotic environments. Persistence in prebiotic systems is not a purely scalar function of energy input, but a spatially and entropically modulated process, one that must be understood in the context of both energy gradients and structural boundary conditions.
5 Thermodynamic Abiogenesis Likelihood Model (TALM)
The Thermodynamic Abiogenesis Likelihood Model (TALM) extends the persistence framework developed in this paper by applying it to planetary-scale and compartmentalized chemical environments. TALM incorporates recent refinements that account for entropic transformation, diffusion constraints, and spatial organization as modulators of molecular persistence.
Entropy-Driven Exploration of Chemical Space
TALM builds on the view that entropy in open, far-from-equilibrium systems is not merely a measure of disorder, but a driver of diversity. Environmental energy fluxes, such as wet-dry cycles or thermal gradients, increase global entropy while enabling localized reductions, forming transient structures and reaction networks.
This entropic dynamism expands the chemical search space, allowing stochastic configurations that, under energy fluctuation, can be filtered via the reaction viability inequality. Thus, entropy facilitates exploration, while thermodynamics imposes a selective constraint.
Reaction Viability in Structured Environments
TALM integrates the augmented viability function:
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t)$ (4)
Here we introduce 2 scaling constants tailored to the environment:
  • $y(t)$ is the net energy balance of a reaction over time
  • $R_{n}$ is a resilience factor, representing structural or contextual redundancy
  • $\Phi(t)$ is the entropic-diffusive penalty, capturing spatial and entropy constraints
We introduce a further 2 modifiers or scaling constants in order to allow us to provide a method of weighting the entropic cost and diffusion loss:
  • $\alpha$ is the scaling constant for $1/\Delta S_{net}$ the thermodynamic cost of maintaining local order
  • $\beta$ is the scaling constant for $D(x,t)$ the diffusion penalty term factor
$\Phi(t)=\alpha\cdot\frac{1}{\Delta S_{net}}+\beta\cdot D(x,t)$
This function enables simulation and comparative evaluation across planetary contexts and captures the fact that in highly diffusive or disordered environments, persistence is penalized unless offset by compartmentalization or external entropy production.
Likelihood Formulation
We model the likelihood of at least one viable, persistent reaction configuration as:
$P_{\text{selective}}=1-(1-P(y^{\prime}(t)\geq 0))^{N}$ (5)
Where:
  • $P(y^{\prime}(t)\geq 0)$ is the probability of a reaction being viable after entropic and spatial correction
  • $N$ is the number of sampled reaction permutations in a given environment
Comparative Context: Earth and Icy Moons
To illustrate the applied potential of the TALM framework, we consider its implications for three planetary environments: early Earth, Europa, and Enceladus. Each of these contexts exhibits the environmental cycling, chemical diversity, and energy gradients necessary to support persistence-aligned reaction dynamics.
Early Earth featured diurnal heating, volcanic geothermal activity, and frequent wet-dry cycles in land-based hydrothermal fields (Deamer, 2011). These conditions likely produced spatial confinement (e.g., in clay layers, mineral pores), low diffusion rates, and fluctuating energy inputs, parameters which, within the TALM framework, increase the probability of $y^{\prime}(t)\geq 0$.
Europa, a Jovian moon, exhibits subsurface oceanic activity influenced by tidal flexing, potentially creating hydrothermal vent systems beneath an icy shell. While lacking sunlight and surface cycling, Europa’s sustained geothermal energy and possible mineral-laden ocean floor could support moderate $P_{\text{selective}}$ values, especially in enclosed or porous mineral environments.
Enceladus, with its plumes of organic-rich water vapor and evidence of internal heating, may also present transient microenvironments with localized gradients and structured surfaces. The TALM framework suggests that even episodic confinement and brief persistence windows could favor molecular configurations that align thermodynamically with available energy pulses.
This comparative perspective demonstrates how TALM can function not just as a theoretical model, but as a tool for ranking or assessing the abiogenesis potential of real-world planetary bodies.
By grounding the concept of reaction viability in energy, entropy, and structure, TALM offers a physically testable and computationally tractable framework for estimating the likelihood of abiogenesis under varying planetary conditions. It enables the comparison of chemical ecosystems by their propensity to sustain persistent molecular configurations, not through genetic inheritance, but through thermodynamic and spatial coherence.
Urability as a Prebiotic Filter
Just as habitability describes environments where life can be sustained, urability refers specifically to environments capable of supporting the origin of life. It characterizes the prebiotic window in which physical and chemical conditions–such as environmental cycling, availability of amphiphiles, and energetically favorable reaction pathways–enable the emergence of molecular complexity prior to heredity or replication Deamer (2021).
In the context of the TALM framework, urability serves as a useful pre–filter for determining which environments warrant thermodynamic modeling. While TALM evaluates whether specific reaction chains can persist under dynamic energy conditions, it assumes that basic prebiotic conditions are met. Urability helps identify planetary settings where those preconditions, such as fluctuating energy, compartmentalization, and chemical substrates, are plausible.
By combining TALM’s probabilistic approach with urability’s environmental screening criteria, we gain a two-tiered method for evaluating the likelihood of abiogenesis: first by assessing whether a planetary body is urable, and second by calculating the statistical probability of persistent reaction networks under its conditions.
6 Illustration of the Model
Reaction chain model
Figure 1 illustrates the logic of the reaction chain model under fluctuating environmental energy.
[…]
Figure 1: Flowchart of the reaction chain viability model.
Energy dynamics over time
Figure 2 illustrates energy dynamics over time, showing the interaction between stored energy and energy used in reactions..
Figure 2: Line chart showing energy dynamics over time.
Numerical Example: Reaction Chain Viability
Reaction Chain Parameters
We define the following reactions:
Reaction $x_{i}$ (Required) $r_{i}$ (Released) $s_{i}$ (Stored)
$R_{1}$ 8 2 1
$R_{2}$ 5 3 2
$R_{3}$ 7 1 1
Case 1: Sufficient Environmental Energy ($z(t)=10$)
  • $R_{1}$: $z=10\rightarrow z-x_{1}+r_{1}=10-8+2=4\rightarrow$ Store $s_{1}=1$
  • $R_{2}$: $z=4\rightarrow z-x_{2}+r_{2}=4-5+3=2\rightarrow$ Store $s_{2}=2$
Result: System collapses at step 3.
Case 2: Higher Energy Input ($z(t)=15$)
  • $R_{1}$: $z=15\rightarrow 15-8+2=9\rightarrow$ Store $s_{1}=1$
  • $R_{2}$: $z=9\rightarrow 9-5+3=7\rightarrow$ Store $s_{2}=2$
  • $R_{3}$: $z=7\rightarrow 7-7+1=1\rightarrow$ Store $s_{3}=1$
Illustrative Example: Thermal Cycling on Early Earth
To illustrate the applicability of the TALM framework under real environmental conditions, consider empirical measurements by Deamer and Damer (2020), who recorded daily temperature fluctuations ranging from 30–90°C in surface hydrothermal pools at Kamchatka. These fluctuations correspond to molecular-scale energy variations on the order of $10^{-20}$ to $10^{-18}$ joules per molecule. Such settings provide ideal testbeds for modeling $z(t)$, the time-dependent environmental energy input. This experimental observation is consistent with the results of Deamer and colleagues, who demonstrated that dehydration–rehydration cycles in hydrothermal fields drive non-enzymatic polymerization of nucleotide-like molecules under prebiotic conditions (DeGuzman et al, 2019) (Ross D & Deamer D, 2019).
$z(t)=z_{\max}\cdot\sin(\omega t)+z_{\text{offset}}$
During daytime, energy peaks may support energetically demanding reactions ($\sum x_{i}$ is high), while nighttime introduces an energy deficit. Only systems with sufficient internal storage $y(t)$, or inherently low energy requirements $x_{i}$, will remain viable during low-input phases.
For example, let a reaction system require $x_{i}=5$ energy units to maintain structural viability and release $r_{i}=1$. If $z(t)$ peaks at 6 and falls to 1 at night, then persistence through a full cycle is only possible for systems that can retain at least $y(t)\geq 3$ units of stored energy from the high-input phase.
This scenario demonstrates TALM’s central principle: differential energy storage enables selective retention. Systems capable of buffering energy during abundance become preferentially stable across environmental fluctuations. Thermal cycling, when combined with compartmentalization and diffusion-limiting substrates, thus constitutes a plausible natural engine for prebiotic selection.
Figure 3: Simulated thermal cycling and entropy export in early Earth conditions. The solid curve represents energy input $z(t)$ over a 48-hour period, modeled as a sinusoidal day–night cycle with baseline energy. The dashed curve shows entropy exported to the environment, $\Delta S_{env}(t)$, phase-shifted to reflect thermal dissipation lag. The shaded region indicates the viable persistence zone, where energy input exceeds the threshold required for reaction continuity. This illustrates how persistent systems must not only align with fluctuating $z(t)$, but also export sufficient entropy to remain thermodynamically viable.
This simplified scenario illustrates how differential energy storage enables selection: systems that store surplus energy are more likely to persist across fluctuating $z(t)$, favoring energy-aligned stability.
Furthermore, such persistence would only be thermodynamically viable if local decreases in entropy, associated with maintaining molecular structure across cycles, are offset by corresponding entropy export to the environment. In this way, day–night thermal cycling could support transient self-organization that satisfies the second law, especially in compartmentalized or surface-adhered systems where diffusion is constrained.
Applying such a model could be used to provide a new approach to the prediction of where life could originate and could be adapted to apply to hydrothermal vents, Europa’s subsurface ocean, or Titan’s hydrocarbon-rich atmosphere, or anywhere where time-variable or localized energy gradients are abundant.
Illustrative Example: Other Planetary Scenarios
In addition to early Earth, and to illustrate the flexibility of the TALM framework, we evaluate five different planetary environments. Each offers different combinations of energy cycling, chemical complexity, and structural confinement whereby we estimate the relative likelihood of life-like persistence using approximate values for energy fluctuation ($P_{E}$), chemical diversity ($P_{C}$), persistence potential ($P_{P}$), and stabilization probability ($P_{S}$). These values are illustrative and intended to demonstrate the framework’s versatility.
1. Rocky Planet with Surface Water (e.g., Early Earth)
Justification: Abundant energy sources, water as solvent, and complex chemistry.
  • $P_{E}\approx 0.9$, $P_{C}\approx 0.8$, $P_{P}\approx 0.3$, $P_{S}\approx 0.7$
$P_{life}\approx 0.1512$
2. Gas Giant Atmosphere (e.g., Jupiter)
Justification: High chemical complexity but turbulent, unstructured environment.
  • $P_{E}\approx 0.95$, $P_{C}\approx 0.7$, $P_{P}\approx 0.05$, $P_{S}\approx 0.4$
$P_{life}\approx 0.0133$
3. Icy Moon with Subsurface Ocean (e.g., Europa)
Justification: Protected liquid ocean, chemical gradients, tidal heating.
  • $P_{E}\approx 0.7$, $P_{C}\approx 0.75$, $P_{P}\approx 0.2$, $P_{S}\approx 0.6$
$P_{life}\approx 0.063$
4. Tidally Locked Exoplanet (e.g., Proxima b-like)
Justification: Persistent light/dark zones may induce stable energy differentials.
  • $P_{E}\approx 0.6$, $P_{C}\approx 0.6$, $P_{P}\approx 0.1$, $P_{S}\approx 0.5$
$P_{life}\approx 0.018$
5. Volcanically Active Super-Earth (e.g., Hypothetical Kepler Candidate)
Justification: High internal energy flux, thick atmosphere, and geochemical cycling.
  • $P_{E}\approx 0.85$, $P_{C}\approx 0.9$, $P_{P}\approx 0.25$, $P_{S}\approx 0.65$
$P_{life}\approx 0.124$
Comparative Summary
These cases demonstrate that while Earth-like planets yield high persistence probabilities, other bodies, including moons and exoplanets, fall within viable ranges, especially when energy stability and chemical diversity align. This reinforces that persistence-driven prebiotic selection is not limited to Earth-like conditions, but may emerge across a spectrum of planetary contexts where structural and thermodynamic constraints are met.
To further highlight the generalizability of the TALM framework, we summarize the core features of these environments in Table 2.
Table 2: TALM Viability Across Planetary Environments
Environment Energy Cycling Features Diffusion Constraint Estimated TALM Viability
Early Earth Wet-dry, geothermal, diurnal Mineral pores, clays Moderate to low High
Europa Tidal heating, hydrothermal Sub-ice mineral surfaces Moderate Moderate
Enceladus Geothermal, episodic plumes Porous rocky interiors Low (in confined settings) Moderate
Titan Seasonal, cryovolcanism Aerosol layers, subsurface ocean Moderate Low to moderate
Mars Geothermal (ancient), brine cycles Subsurface ice/salts Variable Low (present), moderate (past)
Venusian Clouds Photochemical, convection Aerosol droplets High Low but not negligible
Exoplanets / Exomoons Varied (e.g., tidal, rotational, orbital) Minerals, ice fractures, atmosphere Varied Case-dependent

### v7

@@ 3.5.2 The Reaction Viability Inequality
3.5.3 Energy Storage Dynamics and Future Use (S(t))
In open prebiotic systems, a fraction of the energy released during reactions can be retained in chemical or structural form rather than dissipated. This potential stored energy ($s_{i}$) is not immediately usable unless another reaction releases it. Thus, stored energy acts as a possible buffer or reserve, contributing to a system’s future viability but not necessarily to its current energy budget. To represent this explicitly, a storage variable $S(t)$ is introduced to track the time-dependent accumulation and release of stored energy within the system. The rate of change of stored energy is expressed as
$\frac{dS}{dt}=\sum_{i}s_{i}-U(S,t)$
where $\sum_{i}s_{i}$ is the total energy stored in reaction products and $U(S,t)$ is the energy released or utilized from storage as a function of the current state $S(t)$ and environmental conditions. The total available energy at time $t$ therefore becomes
$y(t)=z(t)+\sum_{i}r_{i}-\sum_{i}x_{i}+S(t)$
which ensures that all sources and sinks of energy are explicitly accounted for over a reaction cycle. This formulation preserves conservation of energy across fluctuating input conditions, allowing the persistence inequality $y^{\prime}(t)\geq 0$ to be interpreted as a full-cycle energy balance in which stored energy may buffer temporary deficits in environmental supply.
Figure 1: Figure 1 illustrates the logic of the reaction chain model under fluctuating environmental energy. The flowchart summarizes the operational logic of the reaction-chain persistence model under fluctuating environmental energy.
3.6 Modulators of Persistence
The persistence inequality describes the basic energetic requirement for the continuation or collapse of a reaction sequence. In realistic prebiotic settings, however, this energy balance is influenced by additional physical factors that alter either the effective energy available to a system or the losses it incurs. These modulators include dimensional constraints on the quantities entering the inequality, the dynamics of energy storage and release, the entropic costs associated with local ordering, spatial and diffusive effects, and the resilience provided by network redundancy. Together, these terms extend the core inequality into a form suitable for analyzing prebiotic systems subject to fluctuating, noisy, or spatially structured environments.
3.6.1 Resilience Contribution ($R_{n}$)
The term $R_{n}$ represents the contribution of network redundancy or structural buffering to the overall viability of a reaction sequence. It accounts for the fact that certain reaction networks can sustain temporary local failures, dispersal events, or energetic shortfalls if alternative pathways or parallel processes remain available. To ensure dimensional coherence with the persistence inequality, $R_{n}$ is defined as
$R_{n}=\lambda_{R}\,\mathcal{R}(G),$
where $\lambda_{R}$ is an energy scale (taken here as $RT$ per mole at a reference temperature) and $\mathcal{R}(G)\!\in\![0,1]$ is a normalized measure of network redundancy derived from the topology $G$. Higher values of $\mathcal{R}(G)$ correspond to reaction networks in which multiple reaction sequences or parallel pathways provide buffering against the loss of individual steps. In this way, $R_{n}$ captures the thermodynamic advantage conferred by structural resilience without altering the basic form of the persistence inequality.
3.6.2 Entropy, Diffusion, and Spatial Organization
Entropy production in open thermodynamic systems enables exploration of chemical configuration space, while spatial structure constrains how this exploration unfolds [11, 12]. Local decreases in entropy associated with ordering or assembly must be compensated by entropy export to the surroundings. We therefore define the net entropy balance as:
$\Delta S_{\text{net}}=\Delta S_{\text{env}}-\Delta S_{\text{local}}.$
Here, we take $\Delta S_{\text{env}}\geq 0$ to denote entropy exported to the environment, and $\Delta S_{\text{local}}\geq 0$ to denote the magnitude of entropy reduction associated with local ordering and thus the feasibility condition $\Delta S_{\text{net}}=\Delta S_{\text{env}}-\Delta S_{\text{local}}>0$ ensures compatibility with the second law - systems that decrease local entropy without adequate environmental compensation collapse even if energy thresholds are otherwise met.
Diffusion $D$ further modulates viability. High diffusivity disperses reactants and products, increasing the likelihood of network collapse, whereas confinement in vesicles, pores, or mineral surfaces lowers effective diffusivity and promotes retention. These qualitative effects motivate the introduction of an entropic-diffusive penalty term in the augmented viability function, developed in the next subsection.
Entropy balance and diffusion can be incorporated without altering the structure of the core inequality. We represent the effects through an entropic–diffusive penalty term $\Phi(t)$, which enters the augmented viability function
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t).$
The penalty term $\Phi(t)$ combines the energetic cost of maintaining local order with the losses associated with molecular dispersal. A unit–consistent representation is
$\Phi(t)=\alpha_{S}T\Delta S_{\mathrm{local}}+\beta_{D}\frac{RT}{D_{0}}D(x,t),$
where $\alpha_{S}$ and $\beta_{D}$ are dimensionless scaling factors, $T\Delta S_{\mathrm{local}}$ represents the energetic cost of local entropy reduction, and $(RT/D_{0})D(x,t)$ provides a diffusion‐scaled energetic loss relative to a reference diffusivity $D_{0}=10^{-9}\,\mathrm{m^{2}\,s^{-1}}$. This ensures that $\Phi(t)$ retains units of energy per mole and remains directly comparable with $y(t)$ and $R_{n}$. More detailed models could replace $\Phi(t)$ with system-specific expressions without altering the role of the viability boundary. In this way, the augmented formulation embeds entropy export, spatial confinement, and structural resilience directly into the viability boundary, extending the thermodynamic analysis to realistic prebiotic environments while preserving the simplicity of the core inequality.
3.7 Thermodynamic Consistency and Gibbs Free Energy Formulation
The persistence inequality is fundamentally an energy budget: a reaction network persists when the free energy supplied by the environment, together with any recoverable stored or structural energy, is sufficient to offset the free-energy costs of reaction steps and the losses associated with entropy reduction and diffusion. This condition can be expressed more compactly and transparently using Gibbs free energy, which naturally incorporates both enthalpic and entropic contributions to reaction viability.
For a discrete reaction cycle $n$, we define the net free-energy change of the system as
$\Delta G_{\mathrm{net},n}=\Delta G_{\mathrm{in},n}+\sum_{i=1}^{k}\Delta G_{x,i}+\sum_{i=1}^{k}\Delta G_{r,i}+\Delta G_{\Phi,n}-\Delta G_{R,n}.$
Each term represents a free-energy contribution per cycle, with the following sign conventions:
  • $\Delta G_{\mathrm{in},n}$ is the effective free energy made available to the system from the environment. For a favourable inflow, $\Delta G_{\mathrm{in},n}<0$.
  • $\Delta G_{x,i}>0$ is the free-energy cost associated with initiating or maintaining reaction $i$.
  • $\Delta G_{r,i}$ represents the free-energy change associated with reaction releases. When a reaction releases usable free energy, $\Delta G_{r,i}<0$.
  • $\Delta G_{\Phi,n}>0$ is the free-energy penalty associated with local entropy reduction and molecular diffusion, derived from the entropic–diffusive term $\Phi(t)$.
  • $\Delta G_{R,n}>0$ is the recoverable free-energy contribution arising from structural resilience or redundancy in the network. This term is subtracted because it offsets net free-energy costs.
Persistence requires that the net free-energy change of the system be non-positive:
$\boxed{\Delta G_{\mathrm{net},n}\leq 0.}$
This expresses the requirement that the system, considered over one complete cycle of environmental driving, either dissipates free energy or at least does not accumulate a positive free-energy deficit. Reaction networks with $\Delta G_{\mathrm{net},n}>0$ cannot maintain the fluxes and entropy production required for persistence and therefore collapse.
Mapping between the free-energy formulation and the original energy-budget notation is straightforward:
$z_{n}\;\leftrightarrow\;-\,\Delta G_{\mathrm{in},n},\qquad x_{i}\;\leftrightarrow\;\Delta G_{x,i},\qquad r_{i}\;\leftrightarrow\;-\,\Delta G_{r,i},\qquad R_{n}\;\leftrightarrow\;\Delta G_{R,n},\qquad\Phi_{n}\;\leftrightarrow\;\Delta G_{\Phi,n}.$
Thus the viability condition $y^{\prime}_{n}\geq 0$ is directly equivalent to the thermodynamically grounded condition $\Delta G_{\mathrm{net},n}\leq 0$. The two formulations differ only in representation: the former uses explicit energy accounting, while the latter embeds the same quantities within the standard Gibbs framework for open, driven chemical systems.
3.7.1 Open-System Energy Balance
The persistence inequality can be viewed as a simplified expression of the first law of thermodynamics for an open system in which energy flows continuously between the environment and the reaction network. At any time $t$, the total energy change of the system can be written in the coarse-grained form
$\frac{dU_{\text{sys}}}{dt}=\dot{Q}_{\text{in}}-\dot{Q}_{\text{out}}+\dot{W}_{\text{in}}-\dot{W}_{\text{out}}+\frac{dS}{dt},$
where $\dot{Q}$ represents heat exchange with the environment, $\dot{W}$ represents non-thermal work terms associated with reaction processes, and $\frac{dS}{dt}$ is the net rate of change in stored free energy $S(t)$ defined earlier in the framework. Here $d$ denotes the differential operator, so $\frac{dS}{dt}$ is the continuous-time analogue of the discrete update of $S(t)$.
Persistence requires that the sum of incoming energy and recoverable stored energy meets or exceeds the outgoing dissipative terms. Expressed at the scale of reaction events, this condition reduces to the viability inequality
$y^{\prime}(t)=z(t)+\sum_{i}r_{i}-\sum_{i}x_{i}+S(t)+R_{n}-\Phi(t)\geq 0,$
linking macroscopic energy balance to the microscopic dynamics of reaction viability. This relationship ensures that the framework remains consistent with thermodynamic constraints while retaining a form simple enough to apply across diverse prebiotic environments.
3.7.2 Dimensional Consistency and Energy Balance
To maintain unit coherence within the persistence inequality, all variables are expressed in standard thermodynamic dimensions. The energy terms $x_{i}$, $r_{i}$, and $s_{i}$ represent energy per reaction event, measured in joules per mole (J·mol-1). The environmental input $z(t)$ and residual energy $y(t)$ share the same units. Entropic terms such as $T\Delta S_{\text{net}}$ are expressed in joules per mole through multiplication of temperature (K) by entropy (J·K-1·mol-1). The diffusion factor $D(x,t)$ has dimensions of m2·s-1, and the entropic-diffusive penalty $\Phi(t)$ retains units of energy per mole to remain directly comparable with $y(t)$.
A simple consistency check can be demonstrated by substituting representative values: if a reaction requires $x_{i}=5\times 10^{-20}$ J·molecule-1, releases $r_{i}=2\times 10^{-20}$ J·molecule-1, and the system experiences an environmental input $z(t)=9\times 10^{-20}$ J·molecule-1, then the residual energy is $y(t)=z(t)+\sum r_{i}-\sum x_{i}=6\times 10^{-20}$ J·molecule-1. The units of all terms resolve to energy, confirming dimensional closure of the reaction viability inequality. This ensures that the model remains physically grounded and that persistence, expressed as $y^{\prime}(t)\geq 0$, is evaluated on an absolute energy basis consistent with open-system thermodynamic accounting.
3.8 Relation to Non-Equilibrium Thermodynamic Formalisms
The persistence framework can be situated within the broader developments of non-equilibrium thermodynamics, from classical entropy production theory to modern stochastic approaches. In the canonical formulation of Prigogine and Nicolis, open systems far from equilibrium are governed by generalized thermodynamic forces $X_{i}$ (e.g., chemical potential gradients) and fluxes $J_{i}$ (reaction rates, diffusion currents), with entropy production given by
$\sigma=\sum_{i}J_{i}X_{i}\geq 0,$ (4)
[7]. Dissipative structures arise when thermodynamic fluxes and forces organize to maintain $\sigma>0$. Under these conditions, local decreases in entropy are possible, provided they are compensated by sufficient entropy export to the environment. In this tradition, our condition $\Delta S_{\text{net}}>0$ is directly compatible: persistence requires that local ordering is offset by sufficient entropy release to the environment. Collapse occurs when this balance fails, leading to vanishing fluxes and extinction of the system.
In this setting, the viability condition $y^{\prime}(t)\geq 0$ can be interpreted as the coarse-grained requirement that a reaction system maintains non-vanishing thermodynamic fluxes, and thus positive entropy production, over the relevant time interval. When $y^{\prime}(t)<0$, the system’s capacity to sustain the flux–force pairs $(J_{i},X_{i})$ diminishes, driving $\sigma\to 0$ as reaction rates collapse and the network loses the ability to export entropy. In this sense, the persistence inequality translates the entropy-production criterion $\sigma>0$ for open, driven systems into a local survival condition for specific prebiotic reaction networks.
Recent developments in stochastic thermodynamics extend this picture to finite systems and fluctuating environments. These approaches quantify how entropy production behaves under stochastic forcing and how transient departures from the second law are statistically constrained. The Jarzynski equality[13] and Crooks fluctuation theorem[14] quantify the probability of transient reductions in entropy. Such events are possible, but they are exponentially suppressed, ensuring that the second law holds overwhelmingly on average.
Seifert’s review [15] unifies these results, establishing that entropy production remains the organizing constraint at Mesoscopic scales. Within this context, the persistence inequality provides a coarse-grained survival criterion: reaction networks persist only if $y^{\prime}(t)\geq 0$ holds with high probability across stochastic trajectories. This situates the persistence filter as an operationalization of fluctuation theorems in the specific domain of prebiotic reaction chains.
Esposito and Van den Broeck [16] and Deffner & Lutz [17], further developed non-equilibrium frameworks for driven systems. Their work emphasizes fluctuation-dissipation relations and establishes efficiency bounds for systems undergoing external driving. Our framework differs in focus: rather than maximizing dissipation or efficiency, it identifies a viability boundary where persistence either succeeds or fails. This makes persistence not a variational outcome but a filter imposed by environmental compatibility.
Finally, thermodynamic accounts of abiogenesis such as Michaelian’s dissipation theory [18, 19] propose that life arose as a mechanism for maximizing photon dissipation. This view highlights dissipation as central, but it focuses on a specific energy channel, unlike the more general energy-structure compatibility emphasized by the persistence inequality. Dissipation is necessary but not sufficient: persistence requires meeting the viability criterion in dynamic environments.
Taken together, these connections show that the persistence framework is grounded in non-equilibrium thermodynamics. It condenses broad principles of dissipation, entropy production, and fluctuation constraints into a falsifiable, system-specific viability inequality.